24.03 Paper on Moral Relativism

Matthew Gray

March 12, 1996

Introduction

This paper will discuss various ways of analyzing, evaluating and defending that evaluation of the moral statement:

C: Extraterrestrials (who are without the slightest concern for human life and happiness) ought to avoid injuring humans.

This will be discussed briefly in the non-relativistic frameworks of nihilism and absolutism and the addressed from the perspective of various cognitivist and non-cognitivist versions of relativism. Specifically, the views of Stevenson (emotivism) and Harman will be discussed. Validity of their arguments and plausibility of their premises for evaluation will be covered. Additionally potential objections and discussion of the appropriateness of relativistic treatment of C will be included.

Trivial Cases of Absolutism and Nihilism

In order to best discuss how a relativist might analyze statement C it is useful to consider how the two other broad categories of morality, absolutism and nihilism would consider it. Both views provide questions that must be considered when considering the relativist view.

An absolutist evaluation of C would certainly claim that it has a truth value, and in fact that truth value is completely independent of the moral frameworks of the ET’s, the person stating C and any other parties that one might invoke. The statement is undeniably either true or false to an absolutist. As will be shown later, the various relativists will take exception to various portions of this evaluation.

A nihilist evaluating the same statement would not only fail to endorse its truth, but would explicitly also deny that it’s opposite is true. In fact, the nihilist would claim that both statement C and any “opposite” moral judgment would both be false. This claim would also, incidentally, be independent of the moral frameworks of any of the participants. This view will most notably contrast with the next view to be addressed, emotivism, which also essentially denies any moral statement the ability to be true, but in a sharply different way.

Non-Cognitivist (Emotivist) Relativism

In determining how a non-cognitivist relativist might interpret C, it is difficult to directly compare the non-cognitivist argument with any of the varieties of cognitivist relativism, absolutism or nihilism. This is because the non-cognitivists, in contrast to the other approaches to moral evaluation, do not allow moral statements to have truth values at all. That is, a no-cognitivist would say that a moral statement such as C is no more true (or false) than the statement “Climb the tree”.

Considering this, it is possible to attempt to evaluate the meaning of statement C from a non-cognitivist standpoint, specifically the emotivist, and even more specifically, from the standpoint of the emotivist Stevenson. Other varieties of non-cognitivism won’t be addressed, but most of the comments that apply to emotivism will apply in some senses to other types of non-cognitivism, such as prescriptivism.

The basic interpretation emotivists have for moral statements is that a statement such as “X is wrong” should really be taken to mean “Boo! to X”. Clearly, this translation has no truth value. Using this rule, an emotivist might claim C is best interpreted as “Boo! to ETs injuring humans”, which seems a perfectly sensical, but non-truth-valued sentence.

Stevenson offers a slightly different angle on emotivism, generically translating “X is good” to “X is considered good”. By this system, C would become something like “ETs injuring humans is considered bad”. By using this alternative emotivist interpretation, Stevenson intends to apply any and all of the cognitivist relativist arguments to non-cognitivism. That is if all instances in cognitivist relativism that refer to something as “good” are changed to “considered good”, it yields Stevenson’s emotivism. Some potential problems with this are discussed later.

Appraiser and Agent Relativism

On the cognitivist side, there are two immediately and obviously contrasting approaches. These are appraiser and agent relativism. Each essentially defines in which moral framework a moral statement should be evaluated to determine its truth.

Pure Agent Relativism

The agent relativist alleges that if A states “B ought to do Y”, then A speaks truly if and only if B should do Y in the context of B’s moral framework. In this case, C would be a false statement because in the moral framework of the extraterrestrials, harming humans is irrelevant.

Pure Appraiser Relativism

Appraiser relativism switches the relevant moral framework to that of the person making the statement. That is, if once again A states “B ought to do Y”, then A speaks truly if and only if B ought to do Y in the context of A’s moral framework. In this case, statement C would be true, assuming the author of the statement (the appraiser) believes C.

Hybrid View: Harman

Harman sees an inconsistency in either of these pure approaches. Harman claims that a person A stating that “B ought to do Y” is not completely sensical even if A believes B should do Y, if A knows that B has no reason to do Y. Equally, he claims that it is odd for A to make the same statement if A does not believe that B ought to do Y, even if A knows that B does have moral reason to do Y. This seems to be reason for throwing out both appraiser and agent relativism, but Harman holds onto them both and finds a hybrid.

Harman merges these by asserting that when A says “B ought to do Y” then A speaks truly if and only if the agent B has reasons to do Y and that the appraiser, A in some senses endorses those reasons, and that even the person to whom the statement is being made endorses those reasons as well.^{1} Essentially Harman seems to claim that a moral statement is true if and only if it is true under both an appraiser relativist view and an agent relativist view, as well as being shared due to endorsement of the reasons behind the action.

Harman would say that statement C is odd, if not outright nonsensical. His reasons for this are that the individual making the statement (presumably a human who doesn’t want to be injured by an ET), and the agent of the statement (the ET) do not share a common moral framework surrounding the issue of whether one ought or ought not to injure humans.

1Harman does make it clear that this sort of rewriting only applies to what he calls “Inner Judgments”, and not to all moral statements. This class of moral statement does, however, include statements of the type used throughout this paper

Validity of Arguments

Now that we have established how the various relativists would analyze such a statement, it becomes reasonable to ask whether their analysis is plausible. From there it becomes possible to consider amending these treatments of moral statements to remedy the problems that are found.

Stevenson

In considering the non-cognitivist point of view, specifically Stevenson’s emotivism, it initially seems reasonable to claim that when a person utters statement C they are not in fact making a relativist moral statement, but in fact stating the ET’s injuring humans is not approved of by them. This interpretation, however becomes a bit peculiar when one admits that there might be someone who disagrees with C. That is, it is unlikely that the objector actually disagrees whether they approve or consider good of the ET’s actions, but yet they disagree. This would seem to make normal arguments over moral issues such as C non-sensical at a minimum. While it may be possible to redefine the exact nature of arguments and the like to yield a self-consistent philosophy, it seems unnatural.

Harman

Harman’s interpretation seems to hold more true to the form of interpretation that a rational and consistent relativist might have. Certainly, there is still the issue of disagreements simply being due to the differing moral frameworks of the individuals making the moral statement. This might reduce simply to appraiser relativism, but Harman at least partially saves himself by adding that the appraiser must only endorse the reasons for which the agent is performing the action. This, then, reduces to agent relativism for cases where the appraiser endorses the agents reasons. So, in fact disagreements over moral statements are in fact disagreements over endorsement of the reasons for doing an action. This seems consistent with arguing behaviors.

Lyons levels an objection against Harman that perhaps Harman’s allegation that statement Cis peculiar is not due the the reasons Harman attributes it to. Lyons argues that perhaps the peculiarity perceived in statement Cis not due to a failure of the appraiser and the ET to share common reasons for not injuring humans, but for convictions about the futility of trying to alter the opinion of someone who is likely uninfluencable, or some other equally plausible source. While it is possible in some cases that this is the case, it seems essentially equally reasonable to allow Harman’s point and have Lyons’ objections be metaissues. That is, if a person, in fact, considers C peculiar because they do not believe there is a point in trying to convince the ETs otherwise, perhaps the real issue is whether one should try to convince the ETs otherwise, in which case the argument is now about a separate moral issue “One ought not try to convince someone of something under futile conditions”. This judgment can then be analyzed in Harman’s manner without loss of consistency. This semantic acrobatics could be taken to be a flaw in the structure of Harman’s thesis, or it could merely be necessary due to its subtlety.

Harman’s argument has another possible weakness however. Consider a modified version of C:

C_{2}: Extraterrestrials (who are without the slightest concern for human life and happiness, but who will experience great pain should they injure a human) ought to avoid injuring humans.

Given C_{2}, Harman’s argument would seem to say that this is odd as well. This is because if it is stated by a human (who believes human life is precious and we should therefore avoid injuring humans) A, then the ETs in fact have reason to avoid injuring humans, but their reasons for doing so are by no means endorsed by the human A, despite their actions being endorsed. The claim that C_{2} seems odd, seems odd. Given ETs who have reasons not to injure humans (save their own pain), and a human (A) who doesn’t want humans to suffer (because life is precious), it seems perfectly natural for A to claim that the ETs ought to avoid injuring humans, despite their differing reasons. By Harman’s thesis, this is equally odd as the original version of C. It seems Harman’s merger of appraiser and agent relativism needs to be done more carefully. Certainly in both cases is seems natural to consider the moral framework of the utterer of C or C_{2}, and certainly in both cases the moral framework of the ETs comes into play. It would also seem odd, though, to only require that their moral frameworks dictate an identical course of action, as that seems to divorce the issue from the moral and make it a wholly practical discussion.

Conclusion

The non-cognitivist view apparently has some fundamental flaws, which may be fixable, but in any case seem to go against the normal sense of how a moral relativist might evaluate moral judgments.

Harman’s approach gains much consistency, but it is unclear whether it correctly attributes the cause of the apparent peculiarity of a statement like C. It does seem reasonable though to carefully mend his argument to solve this issue. Further, there may be some conditions where the endorsement of the agent’s reasons by the appraiser is not necessary, but this opens up another entire set of discussions.

Overall, Harman’s approach seems to be fairly complete and practical, though may have some subtle holes and some oversimplification. The non-cognitivist approach, however, is much more fundamentally flawed.